OpenStudy (anonymous):

Use the quadratic equation y = x^2 + 3x – 10. Calculate the discriminant to determine the number of real roots. A. no solution to the equation B. one real root C. no real roots D. two real roots Can someone explain how I should solve this?

OpenStudy (anonymous):

is that x^2 + 3x - 10?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

use the quadratic formula and your roots should be x = -5 x = 2 (x+5) (x-2)

OpenStudy (anonymous):

I might sound dumb, but could you explain the quadratic formula as well? xD

OpenStudy (anonymous):

I think I remember is as an equation with a^2 at the beginning?

OpenStudy (anonymous):

ax^2*

OpenStudy (anonymous):

Quadratic is x = (-b +/- √(b^2 - 4ac))/(2a) or ax^2 + bx + c = 0

OpenStudy (igreen):

Quadratic Formula: \(x = \Large\frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\)

OpenStudy (anonymous):

So I'll plug in the numbers and make the solution?

OpenStudy (anonymous):

0

OpenStudy (igreen):

For solving expressions in the form of: \(ax^2 + bx + c\)

OpenStudy (anonymous):

Plug in x^2 as a + 3x as b - 10 as c

OpenStudy (igreen):

So in this case: a = 1 b = 3 c = -10

OpenStudy (anonymous):

Ok

OpenStudy (igreen):

You will plug them into the quadratic formula and solve.

OpenStudy (anonymous):

What kind of calculator do you have? Is it a graphing calculator?

OpenStudy (anonymous):